# An Empirical Study of Catastrophic Forgetting in Large Language Models During Continual Fine-tuning

### v3

@@ 5.4 Effect of General Instruction Tuning
Figure 6: The performance of general knowledge of the LLAMA-7b model trained on the instruction data and the mixed data.
@@
To further demonstrate the effect of general instruction tuning, we mix 10,000 general instruction data samples from ALPACA (Taori et al., 2023) with the continual instruction tasks to train the LLAMA-7b model. For the sake of brevity, we present the performance of one data split from each evaluation set (MMLU-human, Hellaswag, and Race-middle) to illustrate the effect in Figure 6. The results clearly show that the forgetting during continual instruction tuning can be mitigated to a certain extent by incorporating general instruction data. For instance, the performance of MMLU-human in the initial LLAMA-7b model is 34.72%, but it decreases to 26.8% when trained solely on the instruction data. However, when trained on the mixed data, the performance becomes 30%. These findings further further show that general instruction tuning can help alleviate the CF problem encountered during continual instruction tuning.
@@ 6 Conclusion
In this study, we conducted an empirical investigation into the catastrophic forgetting (CF) phenomenon experienced by large language models (LLMs) during continual instruction tuning. Our findings revealed that the CF problem is generally prevalent in the continual fine-tuning of various LLMs. Moreover, as the model scale increases, LLMs exhibit a more severe degree of forgetting in domain knowledge, reasoning abilities, and reading comprehension skills. Furthermore, our comparative analysis showed that the decoder-only model, BLOOMZ, demonstrates a superior ability to retain knowledge and skills during continual fine-tuning when compared to the encoder-decoder model, mT0. Additionally, we discovered that employing general instruction tuning techniques may help alleviate the CF problem in LLMs. Our empirical study suggests that exploring more effective methods to mitigate CF in LLMs during continual fine-tuning is a promising research direction. When applying LLMs, practitioners should remain vigilant and pay close attention to the issue of knowledge forgetting that may occur after instruction tuning. Addressing this challenge is crucial to ensure the reliable and consistent performance of LLMs in real-world applications.
@@ Appendix A Instruction Task Details
Appendix B Instruction Tuning Results
Here we also show the results of the instruction tuning during continual instruction tuning (Table 7). We mainly present the initial performance and the performance after tuning on the corresponding task to demonstrate the effectiveness of instruction tuning. We use the metrics to measure the model performance on each task following Scialom et al. (2022). In particular, we use SARI for Simp, BLUESore for Emdg, InqQA, Exp, and Rouge-1 for HGen. For example, the initial BLOOMZ-7.1b achieves 51.47% BlueScore in the Exp task, but after continual tuning the model on Simp, Emdg, InqQG, and Exp, the model achieves 68.71% on the Exp task.
Appendix C General Performance Evolution during Instruction Tuning
The detailed results of BLOOMZ-3b, and mT0-3.7b during continual instruction tuning are shown in Figure 8 and Figure 9.
Simp (SARI) Emdg (BS) InqQG (BS) Exp (BS) HGen (R1)
mT0-3.7b 39.01 39.92 48.29 51.70 52.66 56.25 51.62 61.91 30.50 31.77
BLOOMZ-3b 37.95 46.72 46.10 53.27 49.06 59.72 49.75 67.10 27.88 31.72
LLAMA-7b 43.02 46.92 44.28 49.54 43.90 47.54 50.72 54.22 32.42 33.80
ALPACA-7b 45.37 48.22 52.56 54.70 56.91 62.13 52.66 70.49 32.06 36.73
Figure 8: The evaluation of detailed performance in BLOOMZ-3b during continual instruction tuning.
Figure 9: The evaluation of detailed performance in mT0-3.7b during continual instruction tuning.

### v4

@@ 5.4 Effect of General Instruction Tuning
Figure 6: The performance of general knowledge of the BLOOMZ-7.1b and LLAMA-7b model trained on the instruction data and the mixed data. The dashed lines refers to the performance of BLOOMZ-7.1b and LLAMA-7B and the solid ones refer to those of mixed-instruction trained models.
To further demonstrate the effect of general instruction tuning, we mix 10,000 general instruction data samples from ALPACA (Taori et al., 2023) with the continual instruction tasks to train the BLOOMZ-7b and LLAMA-7b model. For the sake of brevity, we present the performance of one data split from each evaluation set (MMLU-human, Hellaswag, and Race-middle) to illustrate the effect in Figure 6. The results clearly show that the forgetting during continual instruction tuning can be mitigated to a certain extent by incorporating general instruction data. For instance, the performance of MMLU-human in the initial LLAMA-7b model is 34.72%, but it decreases to 26.8% when trained solely on the instruction data. However, when trained on the mixed data, the performance becomes 30%. These findings further further show that general instruction tuning can help alleviate the CF problem encountered during continual instruction tuning.
@@ 6 Conclusion
In this study, we conducted an empirical investigation into the catastrophic forgetting (CF) phenomenon experienced by large language models (LLMs) during continual instruction tuning. Our findings revealed that the CF problem is generally prevalent in the continual fine-tuning of various LLMs. Moreover, as the model scale increases, LLMs exhibit a more severe degree of forgetting in domain knowledge, reasoning abilities, and reading comprehension skills. Furthermore, our comparative analysis showed that the decoder-only model, BLOOMZ, demonstrates a superior ability to retain knowledge and skills during continual fine-tuning when compared to the encoder-decoder model, mT0. Additionally, we discovered that employing general instruction tuning techniques may help alleviate the CF problem in LLMs. Our empirical study suggests that exploring more effective methods to mitigate CF in LLMs during continual fine-tuning is a promising research direction. Meanwhile, since our work is an empirical study and is constrained by the computation resources, there is still large room to investigate the forgetting phenomenon such as a larger model scale (70b or larger) since a larger model may need less parameter changes to fit the downstream tasks. When applying LLMs, practitioners should remain vigilant and pay close attention to the issue of knowledge forgetting that may occur after instruction tuning. Addressing this challenge is crucial to ensure the reliable and consistent performance of LLMs in real-world applications.